# Unified Modeling Language Code Generation from Diagram Images Using Multimodal Large Language Models

### v1

@@ 4.1.1 Training Trends
The computational demands of machine learning model training depend on the size of the dataset, the complexity of the dataset, and the model architecture. Tables 3, 4, 5, and 6 detail the metrics for training time and computational performance across activity and sequence datasets of varying sizes (small to extra-large) for both the 7B and 13B model architectures. While activity and sequence datasets share trends in step rates and throughput, differences in complexity lead to notable disparities in computational demands. Step rates decrease slightly as the dataset size grows. For instance, the 7B full model processes 0.081, 0.080, and 0.079 steps per second for the small, medium, and large activity datasets. Small reductions in step rates reflect the increasing computational demands of larger datasets while demonstrating the model’s consistency. Similarly, the 7B full model maintains a steady 0.108 steps per second across small, medium, and large sequence datasets, highlighting the lighter computational burden of sequence data. These results confirm the robustness of the model architecture in maintaining efficiency across different scales and types of datasets.
@@ 4.3 Result Overview
LoRA fine-tuning yielded results comparable to full fine-tuning but revealed trade-offs related to hardware constraints. While LoRA required fewer GPUs (two versus four), it experienced longer training durations and higher FLOS due to the smaller batch size (8 for LoRA compared to 16 for full fine-tuning). This configuration limited LoRA’s training throughput, though it often outperformed fully fine-tuned models on metrics such as BLEU and SSIM, particularly for complex activity diagrams. LoRA’s targeted parameter updates captured UML structure and syntax, underscoring its utility in resource-constrained scenarios. Despite its longer training times, these findings highlight LoRA’s potential as a practical approach for tasks requiring high fidelity in structured data. Overall, the results validated the hypotheses that larger models and datasets enhance performance and that fine-tuning, particularly LoRA, significantly improves model scalability and accuracy. While LoRA fine-tuning exhibited slower overall training times, its performance remained strong across all metrics, emphasizing its adaptability and efficiency in structured data tasks. Future work can further refine UML code generation models by leveraging larger datasets and exploring advanced alignment techniques.
@@ 5 Discussion
This study investigates the role of the LLaVA-1.5 model in automating UML diagram-to-code generation, expanding on prior work by Conrardy and Cabot. While their research prioritizes syntactic precision and prompt design for models like GPT-4 and CogVLM [7, 30], our broader approach asses both structural fidelity and computational efficiency. By incorporating BLEU and SSIM metrics, we evaluate textual and visual accuracy, offering a more comprehensive performance assessment across various datasets and model configurations. Results emphasize the critical role of dataset scale and fine-tuning strategies. Larger LoRA-tuned models, particularly the 13B configurations, consistently outperform baseline models by achieving higher BLEU and SSIM scores. These models excel at generating structurally accurate UML diagrams, effectively reducing syntax errors and mismatches. Fine-tuning, especially with LoRA, demonstrates how targeted parameter updates enhance performance without overwhelming computational costs, making advanced models accessible to resource-limited environments.
Compared to Conrardy and Cabot’s focus on prompt refinement, our study addresses computational scalability, a key factor in adoption. While they achieved impressive results with structural accuracy for class diagrams, we highlight the importance of scalability and model adaptability for complex behavioral diagrams like sequence and activity workflows. Both studies demonstrate the growing potential of multimodal models in software engineering. Still, our inclusion of error analysis—quantifying mismatches and syntax issues—offers deeper insights into areas where models can improve further.
Our research also reveals that model misinterpretations, such as defaulting to incorrect diagram types, are more prevalent in smaller datasets. However, fine-tuned models correct these issues by adhering more closely to prompts and minimizing mismatches. These findings demonstrate that larger datasets and well-tuned models enable MM-LLMs to handle dynamic workflows effectively. By addressing key challenges such as structural errors, prompt misinterpretations, and computational bottlenecks, our work sets the stage for scalable and efficient UML diagram automation.
@@ 5.1 Avenues for Further Study
The BLEU metric has been a foundational tool for evaluating code generation; however, its reliance on syntactic $n$-gram matching may limit its applicability for UML tasks, where visual fidelity and structural relationships are also critical. CodeBLEU extends BLEU by incorporating abstract syntax tree (AST) analysis and data flow consistency, aligning evaluations more closely with functional correctness. However, its current design focuses on traditional programming languages such as Python or C [26]. Exploring how similar techniques might be adapted for UML tasks could open avenues for aligning evaluations with textual and visual aspects. Furthermore, approaches such as pass@k, which evaluates correctness across multiple outputs, and frameworks such as EvalPlus, which uses mutation-based testing to identify potential model vulnerabilities [16], present interesting opportunities for advancing UML evaluation methods by integrating textual and diagrammatic assessments. Emerging multimodal models such as LLaVA-NeXT [14] and LLaVA-OneVision [12], which incorporate advanced OCR, multi-resolution processing, and cross-modal knowledge transfer, offer promising directions for improving UML automation by better handling various types of diagram and maintaining accuracy at varying resolutions. Similarly, expanding datasets to include a broader range of UML types, such as class diagrams [28], and leveraging synthetic and real-world examples, could enhance model generalization and robustness while reducing overfitting [34]. These avenues highlight the potential for future exploration into developing efficient, scalable, and broadly applicable UML automation tools that balance computational efficiency with usability.
@@ 6 Conclusion
The study highlights opportunities for future exploration, such as modifying advanced metrics like CodeBLEU and pass@k, which offer deeper insights into functional correctness, and leveraging state-of-the-art models like LLaVA-NeXT and LLaVA-OneVision to enhance reasoning and scalability. Expanding datasets to include additional UML types, such as class diagrams, and incorporating synthetic and real-world examples will strengthen model generalization and robustness. These advancements will refine UML automation tools that align with the evolving requirements of modern software engineering workflows.
Our contributions demonstrate the potential of intelligent, scalable AI systems to automate complex design tasks and bridge theoretical advances with practical applications. The framework offers software engineering solutions and broader fields requiring structured and dynamic representations, including healthcare, education, and industrial automation. Establishing a strong foundation for AI-driven design and implementation creates new opportunities for innovation across diverse industries, advancing intelligent system integration and automation.
@@ Acknowledgments
The authors thank the University of Oklahoma’s Supercomputing Center for Education & Research for providing access to their supercomputing resources, including the Schooner cluster. We also thank the Data Institute for Societal Challenges for facilitating GPU resources through OSCER’s infrastructure. In addition, we acknowledge AFWERX for their support and contributions. Finally, the authors extend their deepest gratitude to Dr. Andrew H. Fagg for his invaluable feedback on the thesis that formed the foundation of this work. His guidance were helped shape the direction and quality of the research.
@@
This project was funded by the U.S. National Science Foundation under the Accelerating Research Translation program 2331409 and the U.S. Air Force Small Business Innovation Research program.
@@
Supplementary Figures
Additional supplementary figures are available in the supplementary material file.
@@
Fig. S4: Ground truth sequence UML diagram illustrating the original reference structure
Fig. S5: Generated UML diagram with high SSIM and BLEU scores, showing strong structural and textual alignment with the ground truth.
Fig. S6: Ground truth activity UML diagram illustrating the original reference structure.
Fig. S7: Generated UML diagram with low SSIM and BLEU scores, highlighting structural misalignments and textual discrepancies compared to the ground truth.

### v2

@@ 3.1 Model Performance
3.1.1 Real-World Evaluation and Limitations
Fig. 14: Model performance on the real-world activity diagram dataset, showing BLEU and SSIM scores across baseline, full fine-tuning, and LoRA configurations with the small dataset models. Compared to synthetic data, improvements from fine-tuning are more modest.
Fig. 15: Model performance on the real-world sequence diagram dataset. Fine-tuned models, for the small dataset models, show only minor gains over the baseline, suggesting that visual mismatches between training and test data affect generalization.
Fig. 16: Comparison of BLEU and SSIM scores across real-world activity and sequence diagrams. Results indicate that both full and LoRA fine-tuning yield limited improvements over the baseline, underscoring the importance of aligning synthetic training data with real-world visual styles.
Performance on real-world diagrams remained limited, as shown in Figures 14, 15, and 16. While increasing model size and training duration led to slight gains—BLEU scores for the activity test set improved from 0.012 of the orignial 7B model to 0.038 for the extra large LoRA training data model, and SSIM increased from 0.213 to 0.643—these improvements were modest. Results showed an even smaller increase in sequence data from the small training models to the extra large training models. The relatively small improvements underscore the gap between training conditions and real-world application.
Supplementary figures, Fig. S8 through Fig. S11, visualize these limitations, contrasting model-generated outputs with actual UML diagrams. The differences in terminology, formatting, and syntactic structure make it clear that synthetic data lacks realism in key areas. These observations point to the need for more representative training data that better mirrors real-world patterns. Given the difficulty of scaling real-world data collection, a practical path forward involves refining synthetic data generation to better simulate real-world use. Designing datasets around specific application domains—such as authentication, e-commerce, or user management—and adopting realistic naming conventions could significantly enhance generalization.
@@ 3.2.1 Training Trends
The computational demands of machine learning model training depend on the size of the dataset, the complexity of the dataset, and the model architecture. Tables 4, 5, 6, and 7 detail the metrics for training time and computational performance across activity and sequence datasets of varying sizes (small to extra-large) for both the 7B and 13B model architectures. While activity and sequence datasets share trends in step rates and throughput, differences in complexity lead to notable disparities in computational demands. Step rates decrease slightly as the dataset size grows. For instance, the 7B full model processes 0.081, 0.080, and 0.079 steps per second for the small, medium, and large activity datasets. Small reductions in step rates reflect the increasing computational demands of larger datasets while demonstrating the model’s consistency. Similarly, the 7B full model maintains a steady 0.108 steps per second across small, medium, and large sequence datasets, highlighting the lighter computational burden of sequence data. These results confirm the robustness of the model architecture in maintaining efficiency across different scales and types of datasets.
@@ 3.4 Result Overview
LoRA fine-tuning yielded results comparable to full fine-tuning but revealed trade-offs related to hardware constraints. While LoRA fine-tuning significantly reduces the number of trainable parameters and lowers memory consumption, our findings highlight an important trade-off: increased training time and higher total FLOPs compared to full fine-tuning. Despite requiring fewer GPUs (two versus four for full fine-tuning), LoRA exhibited slower step rates and longer convergence times. This was largely due to hardware and batch size constraints. The LLaVA training pipeline is designed to maintain a consistent global batch size. To match the original global batch size of 128 (as used in the official 8-GPU setup), we configured LoRA training on two GPUs with a per-device batch size of 16 and four accumulation steps. In contrast, standard fine-tuning used four GPUs with a batch size of 8 per device and the same accumulation steps. While the global batch size was preserved, using fewer GPUs with larger per-device batches increased memory pressure and reduced per-step efficiency. For example, across the 13B model on extra-large activity datasets, LoRA fine-tuning consumed approximately $2.772\times 10^{15}$ FLOS over 9 hours, compared to $2.48\times 10^{15}$ FLOS in 5.95 hours for full fine-tuning. This suggests that LoRA’s efficiency gains in memory and GPU requirements come at the cost of computational throughput—likely due to differences in batch size and limited GPU resources. However, the flexibility it offers—especially for resource-constrained setups—often outweighs the extended runtime, making it a valuable strategy for scalable domain adaptation.
Although real-world data was not used during training, we evaluated model generalization on a small set of naturally sourced UML diagrams. The results, while showing small improvements with training, revealed a notable performance gap compared to synthetic test data. The discrepancy is likely due to differences in naming structure and semantic coherence, as the synthetic data was generated using randomized labels. Nonetheless, the experiment demonstrates the feasibility of using synthetic data for this task and highlights opportunities for improving generalization. The extensive synthetic datasets reveal that, regardless of other factors, with sufficient data the models can effectively learn to represent UML code from images.
@@ 4 Discussion
This study investigates the role of the LLaVA-1.5 model in automating UML diagram-to-code generation, expanding on prior work by Conrardy and Cabot. While their research prioritizes syntactic precision and prompt design for models like GPT-4 and CogVLM (Conrardy and Cabot, 2024; Wang et al., 2024), our broader approach asses both structural fidelity and computational efficiency. By incorporating BLEU and SSIM metrics, we evaluate textual and visual accuracy, offering a more comprehensive performance assessment across various datasets and model configurations. Results emphasize the critical role of dataset scale and fine-tuning strategies. Larger LoRA-tuned models, particularly the 13B configurations, consistently outperform baseline models by achieving higher BLEU and SSIM scores. These models excel at generating structurally accurate UML diagrams, effectively reducing syntax errors and mismatches. Fine-tuning, especially with LoRA, demonstrates how targeted parameter updates enhance performance without overwhelming computational costs, making advanced models accessible to resource-limited environments.
Compared to Conrardy and Cabot’s focus on prompt refinement, our study addresses computational scalability, a key factor in adoption. While they achieved impressive results with structural accuracy for class diagrams, we highlight the importance of scalability and model adaptability for complex behavioral diagrams like sequence and activity workflows. Both studies demonstrate the growing potential of multimodal models in software engineering. Still, our inclusion of error analysis—quantifying mismatches and syntax issues—offers deeper insights into areas where models can improve further. Our research also reveals that model misinterpretations, such as defaulting to incorrect diagram types, are more prevalent in smaller datasets. However, fine-tuned models correct these issues by adhering more closely to prompts and minimizing mismatches. These findings demonstrate that larger datasets and well-tuned models enable MM-LLMs to handle dynamic workflows effectively. By addressing key challenges such as structural errors, prompt misinterpretations, and computational bottlenecks, our work sets the stage for scalable and efficient UML diagram automation.
@@ 4.1 Avenues for Further Study
The BLEU metric has been a foundational tool for evaluating code generation; however, its reliance on syntactic $n$-gram matching may limit its applicability for UML tasks, where visual fidelity and structural relationships are also critical. CodeBLEU extends BLEU by incorporating abstract syntax tree analysis and data flow consistency, aligning evaluations more closely with functional correctness. However, its current design focuses on traditional programming languages such as Python or C (Ren et al., 2020). Exploring how similar techniques might be adapted for UML tasks could open avenues for aligning evaluations with textual and visual aspects. Furthermore, approaches such as pass@k, which evaluates correctness across multiple outputs, and frameworks such as EvalPlus, which uses mutation-based testing to identify potential model vulnerabilities (Liu et al., 2023), present interesting opportunities for advancing UML evaluation methods by integrating textual and diagrammatic assessments.
Emerging multimodal models such as LLaVA-NeXT (Liu et al., 2024b) and LLaVA-OneVision (Li et al., 2024), which incorporate advanced OCR, multi-resolution processing, and cross-modal knowledge transfer, offer promising directions for improving UML automation by better handling various types of diagram and maintaining accuracy at varying resolutions. While our experiments centered on LLaVA-based models, we acknowledge the importance of incorporating additional benchmark models to broaden the comparative landscape. Our primary objective was to demonstrate the feasibility of using a multimodal LLM for large-scale UML code generation. Within this scope, LLaVA-1.5 at both 7B and 13B scales proved sufficient for validating our approach. Nevertheless, including models such as Deepseek R1 (DeepSeek-AI et al., 2025), Claude, or CogVLM in future work would offer a more comprehensive perspective on architectural differences, visual reasoning capabilities, and generalization.
Our models achieved strong performance on synthetic datasets, gains on more realistic UML samples were notably more modest. As illustrated in Figs. 14 and 15, fine-tuned models demonstrated only marginal improvements over the baseline when evaluated on real-world diagrams. This discrepancy stems from synthetic diagrams’ structure, which fail to fully capture the artifacts commonly found in typical UML diagrams. To bridge the realism gap, future iterations of our synthetic dataset could categorize diagrams into common application domains such as banking, e-commerce, authentication workflows, and web development. The thematic structuring would align synthetic diagrams more closely with real-world use cases, thereby improving their semantic fidelity. Additionally, incorporating domain-specific terminology and realistic naming conventions within these categories would further enhance the authenticity of the synthetic data. Nevertheless, our results strongly suggest that synthetic data can serve as a powerful training signal with sufficient volume and variation: MM-LLMs exhibit a clear capacity to learn to accurately reconstruct UML code from images. The result highlight a promising direction—by further enhancing the visual realism and variability of synthetic training data, future models may bridge the domain gap and achieve high-fidelity UML code generation even in real-world scenarios.
@@ 5 Conclusion
The study highlights opportunities for future exploration, such as modifying advanced metrics like CodeBLEU and pass@k, which offer deeper insights into functional correctness, and leveraging state-of-the-art models like LLaVA-NeXT and LLaVA-OneVision to enhance reasoning and scalability. Expanding datasets to include additional UML types, such as class diagrams, and incorporating synthetic and real-world examples will strengthen model generalization and robustness. These advancements will refine UML automation tools that align with the evolving requirements of modern software engineering workflows. Our contributions demonstrate the potential of intelligent, scalable AI systems to automate complex design tasks and bridge theoretical advances with practical applications. The framework offers software engineering solutions and broader fields requiring structured and dynamic representations, including healthcare, education, and industrial automation. Establishing a strong foundation for AI-driven design and implementation creates new opportunities for innovation across diverse industries, advancing intelligent system integration and automation.
@@ Acknowledgments
The authors thank the University of Oklahoma’s Supercomputing Center for Education & Research for providing access to their supercomputing resources. We also thank the Data Institute for Societal Challenges for facilitating GPU resources through OSCER’s infrastructure. In addition, we acknowledge AFWERX for their support and contributions. Finally, the authors extend their deepest gratitude to Dr. Andrew H. Fagg for his invaluable feedback on the thesis that formed the foundation of this work. His guidance were helped shape the direction and quality of the research.
CRediT authorship contribution statement
Averi Bates: Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Writing - original draft, Writing - review & editing, Visualization. Ryan Vavricka: Data curation, Software, Writing - review & editing. Shane Carleton: Conceptualization, Supervision, Project administration, Funding acquisition. Ruosi Shao: Writing - review & editing. Chongle Pan: Conceptualization, Supervision, Project administration, Writing - review & editing.
@@
This project was funded by the U.S. National Science Foundation under the Accelerating Research Translation program 2331409 and the U.S. Air Force Small Business Innovation Research program. Additional funding for publication was provided by the University of Oklahoma Libraries’ Open Access Fund
@@
Data Availability
The training and testing UML data supporting the findings of this study are available and can be found at https://doi.org/10.5281/zenodo.15103682. Please note that this excludes any associated prompts. Access to the data will be granted for non-commercial research purposes.
Appendix A Supplementary Figures
The following is the supplementary material related to this article.
@@ Appendix A Supplementary Figures
Fig. S4: Ground truth UML diagram used for comparison.
Fig. S5: Generated UML sequence diagram with high SSIM and BLEU scores, showing strong structural and textual alignment with the ground truth.
Fig. S6: Ground truth UML diagram illustrating the original reference structure.
Fig. S7: Generated UML activity diagram with low SSIM and BLEU scores, highlighting structural misalignments and textual discrepancies compared to the ground truth.
Fig. S8: Ground truth UML diagram illustrating the original reference structure of real world data rather than synthetic.
Fig. S9: Example of a generated UML activity diagram with a moderatly high SSIM score and low BLEU score, highlighting structural and syntactic issues that arise from training on unstructured synthetic data.
Fig. S10: Ground truth UML diagram illustrating the original reference structure of real world data rather than synthetic.
Fig. S11: Example of a generated UML sequence diagram with low SSIM and BLEU scores, highlighting structural and syntactic issues that arise from training on unstructured synthetic data.